OpenStudy (ledah):

The graph of the system is 2x+y=4 3x-y=1

OpenStudy (anonymous):

so basiccaly is y = -2x +4 and y = 3x-1 if written in general form. both intersecting lines; have different gradient. so what's the gist of the question? o.o

OpenStudy (ledah):

o_o not sure what you mean by that

OpenStudy (anonymous):

hmm u said a graph of a system is 2x+y=4 then what you meant by 3x-y=1

OpenStudy (ledah):

that was the whole problem it just posted it funny

OpenStudy (anonymous):

LOL

OpenStudy (ledah):

o_o

OpenStudy (anonymous):

I gave you a medal for best answer. but what i really meant was I think its the best answer ever

OpenStudy (ledah):

What answer o_o xD confused.....

OpenStudy (anonymous):

"that was the whole problem..."

OpenStudy (ledah):

It was xD

OpenStudy (anonymous):

do you need to solve for x and y or do you need to graph it

OpenStudy (ledah):

Its a matching exercise. let me see if i can post this link

OpenStudy (ledah):

o_o